OpenStudy (anonymous):

What is the value of k if x3 + kx2 + 7x + 5 is divided by x + 6 and gives a remainder of −1?

OpenStudy (tkhunny):

How's your synthetic division?

OpenStudy (mathmale):

Please express exponentiation properly. It's not x3 + kx2 + 7x + 5, but rather \[x^3+ kx^2+7x+5\]

OpenStudy (mathmale):

Next question is what to do with that last term, the 5. If the remainder is -1, should we add -1 to 5 (to obtain a constant term of 4), or should we subtract -1 from 5 (to obtain a constant term of 6. Personally I have to experiement to answer my own question.

OpenStudy (mathmale):

(x+6) is a factor (given). Thus, we use -6 as the divisor in synthetic division. First I assume that 4 as a constant term would result in no remainder: -6 / 1 k 7 4 -6 -6k+36 36k-258 What value of k would result in a rem. ______________________________________ of zero? 1 k-6 -6k+43

OpenStudy (mathmale):

This test does not necessarily provide the correct value for the unknown, k. Would anyone like to propose a more systematic method that would reveal the correct k value more efficiently?

OpenStudy (tkhunny):

?? 36k - 254 = -1 Solve.

OpenStudy (freckles):

The remainder theorem can be used.

OpenStudy (freckles):

Let P(x) be a polynomial. \[\frac{P(x)}{x-c}=Q(x)+\frac{R}{x-c} \\ P(x)=Q(x) (x-c)+R\] Where R is a constant. \[P(c)=Q(x)(c-c)+R \\ P(c)=Q(x)(0)+R \\ P(c)=0+R \\ P(c)=R\]

OpenStudy (tkhunny):

Yup. Just did. Less formally, with the synthetic division, of course.

OpenStudy (freckles):

I understand that. Someone said something about finding another way too though so that is why I presented this way. Learning this theorem also means less work. Since all you have to do here is solve P(-6)=-1 for k where P is that third degree polynomial given